OpenStudy (anonymous):

The first two terms of an infinite GP ( geometric progression ) are together equal to 5 and every term is 3 times the sum of all the terms that follows it; find the series.

terenzreignz (terenzreignz):

Your sequence should take the form of \[\huge \{ ar^{n }\}\]where a is your initial value and r is your common ratio. Just find a and r.

terenzreignz (terenzreignz):

You know two things, the first two terms add up to 5, so... \[a + ar = 5\] You have two unknown, and one linear equation, maybe the second fact you know would lead to another one. You need two linear equations to get to a definite answer if you have two unknowns.

OpenStudy (anonymous):

that i have already done.. please let me know further

terenzreignz (terenzreignz):

Do you know how to evaluate an infinite geometric series (ie, the sum of the infinite terms of a geometric progression) ?

OpenStudy (anonymous):

\[S_{\infty} = \frac{a}{1-r}\]

terenzreignz (terenzreignz):

Where a is the initial term and r is the common ratio, correct. Now it says that every term is equal to three times the sum of all the terms that follow it. How do we interpret this?

OpenStudy (anonymous):

Sum of an infinite G.P is a multiple of \[\frac{r}{1-r}\] factor with a previous term that means : \[ar^n+ar^{n+1}+ar^{n+2}.....\infty \] = \[ar^{n-1}\frac{r}{1-r}\]

terenzreignz (terenzreignz):

Simply put, for all non-negative integers, n \[\huge ar^{n}=?\]

terenzreignz (terenzreignz):

Well, it says that every term is equal to the sum of the succeeding terms, ie, the rest of the geometric series, so... \[\huge ar^{n}=\sum_{k=1}^{\infty}ar^{n+k}\] Go ahead and simplify this.

OpenStudy (shubhamsrg):

or,,simply 1st term = 3* (sum of all the terms - 1st term) 4a = 3a/(1-r) 2 eqns, 2 variables..